OpenStudy (anonymous):

Use the distributive property to simplify the expression below: −2(d − 4) + 3d

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow -2d + 8 + 3d }\) \(\Large \color{MidnightBlue}{\Rightarrow d + 8 }\)

OpenStudy (anonymous):

Evaluate: 14 − 10 ÷ 2 + 3

Parth (parthkohli):

Use BODMAS in this one. \(\Large \color{MidnightBlue}{\Rightarrow 14 - (10 \div 2) + 3 }\) \(\Large \color{MidnightBlue}{\Rightarrow 14 - 2 + 3 }\) \(\Large \color{MidnightBlue}{\Rightarrow 14 + 1 }\) \(\Large \color{MidnightBlue}{\Rightarrow 15 }\)

OpenStudy (anonymous):

Bryan..you are posting very simple question. May I know ur age and your academic standard pls ? are you in school ?

OpenStudy (anonymous):

Simplify the following expression. 2x − 2y + 3z − 2x − y + 2z = ?

Parth (parthkohli):

Yes, as vikas said

OpenStudy (anonymous):

Are you here to get your homework solved ? They are so easy. You should try it yourself.

OpenStudy (anonymous):

im in school

Parth (parthkohli):

Operate on like terms, Bryan.

OpenStudy (anonymous):

im in 10th grade im actually good in math got high fcat scores but lazy to do work

Parth (parthkohli):

I'm just 12...can't believe you ain't able to do all this :/

OpenStudy (anonymous):

which country you from ? Any of this question should not take more than 2 seconds to solve. I am serious!

OpenStudy (anonymous):

im from usa

OpenStudy (anonymous):

teach me how to do them quick?

Parth (parthkohli):

Operate on like terms, as I said.

OpenStudy (anonymous):

ok im gonna try to do this math problem Simplify the following expression. 2x − 2y + 3z − 2x − y + 2z = ?

OpenStudy (anonymous):

3y+5z?

Parth (parthkohli):

Nope

OpenStudy (anonymous):

-3y?

OpenStudy (anonymous):

2x − 2y + 3z − 2x − y + 2z.. Lets rearrange the like terms as (2x− 2x)+(3z+2z)+(-2y-y) and now do the calculation (0)+(5z)+(-3y) = 5z-3y Got it ?

Parth (parthkohli):

Yes =)

OpenStudy (anonymous):

im pretty sure i got it

OpenStudy (anonymous):

thanks for your help

OpenStudy (anonymous):

mind giving me a problem so i can solve?

OpenStudy (anonymous):

Its Ok..though you need to learn and practice a lot...

Parth (parthkohli):

Solve this \(\Large \color{MidnightBlue}{\Rightarrow 15x^2y + 9xy + 4 - 9x^2y - 3xy - 15 }\)

OpenStudy (anonymous):

36y+6xy-11?

OpenStudy (anonymous):

or 42xy-11?

OpenStudy (anonymous):

did i do it correcT?

Parth (parthkohli):

No!

OpenStudy (anonymous):

what did i do wrong?

OpenStudy (anonymous):

the terms which are of same form..bring them together and bracket them as i did...in ur problem 15x2y+9xy+4−9x2y−3xy−15 = (15x2y-9x2y)+(9xy-3xy)+(4-15)= 6x2y+6xy-11 got it this time ?

OpenStudy (anonymous):

Bryan..u ok ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

just learning

OpenStudy (anonymous):

oh so i basically could of gotten it correct if i hadent solved the 6 to the power

OpenStudy (anonymous):

i mean squared

OpenStudy (anonymous):

vikash you ok?

OpenStudy (anonymous):

The problem is that you are not confident of your method..you are coming up with multiple answers for the same question. You need to understand the mechanism well and apply it uniformly in all other problems.. Let me tell you..whenever you solve some new type of problem..do not rush to the new one immediately. Hold on..have a look at the problem you just solved..understand the concept..see how it applies and only then move to the next problem. If you follow this you will never repeat your mistakes.

OpenStudy (anonymous):

do not rush to the next problem immediately. *

OpenStudy (anonymous):

ok i will trry not to